Question 21

Allen Inc. took out a 1-year, 8%, $120,000 loan on September 1, 2019. Interest is due upon maturity of the loan. The loan and interest must be paid back on August 31, 2020. As of December 31, 2019, what amount, if any, should Allen Inc. report for interest payable?

a- $9,600

b- $0

c- $6,400

d- $3,200

Question 22

An example of a closing entry would be:

a- Closing Sales Revenue to Retained Earnings.

b- Closing Net Income to Common Stock.

c- Closing Accounts Receivable to Cash.

d- Closing Depreciation Expense to Accumulated Depreciation.

Question 23

Herring Corp has a post-closing Retained Earnings balance of $107,400. If during the year Herring paid dividends of $110,000 and had net income of $142,000, then what was Herring Corp’s beginning Retained Earnings balance?

a- $32,000

b- $249,400

c- $139,400

d- $75,400

Question 24

Accountants are responsible for measuring various operating, investing and financing activities. Which of the following correctly matches the activity with its type?

a- Financing – paying dividends to its shareholders

b- Investing – selling additional stock in the company to the investors

c- Investing - paying wages for the month

d- Operating – purchasing equipment to use in operations

Question 25

On September 1, 2019, Fairacres Equipment Leasing, Inc. paid $48,000 for 24 months of advance rent on its warehouse. What would be the amount of prepaid rent on the December 31, 2019 financial statements for Fairacres under accrual-basis accounting?

a- $0

b- $40,000

c- $16,000

d- $8,000
Answer #1

21. Interest payable for 2019

= 120,000 * 8% * 4 months/12 months

= 3200

Option D

22. Option D

All the temporary accounts are classed at the end of accounting period. Depreciation expense being a temporary account, is closed to Accumulated Depreciation account.

23. Ending retained earnings = Beginning retained earnings + Net income - Dividends paid

107,400 = Beginning retained earnings + 142,000 - 110,000

Beginning retained earnings = 75,400

Option D

24. Option A

Selling additional stock is a financing activity. Paying wages is a operating activity. Purchasing equipment is a investing activity.

25. Rent expense for the year 2019 = (48,000/24)*4 months = 8,000

Balance in prepaid rent account on Dec 31,2019

= 48,000 - 8,000

= 40,000

Option B